OpenStudy (anonymous):

@jim_thompson5910 Write the equation of a line in point-slope form that has a slope of -1 and passes through the point (-2, 5) (Points : 2) y + 5 = -1(x – 2) y + 2 = 1(x – 5) y + 1 = 5 (x + 2) y – 5 = -1(x + 2)

jimthompson5910 (jim_thompson5910):

use y - y1 = m(x - x1)

jimthompson5910 (jim_thompson5910):

m is your slope (x1,y1) is the point that the line goes through

OpenStudy (anonymous):

WHAT NUMBERS DO I P[UT IN??

jimthompson5910 (jim_thompson5910):

whats the given slope

OpenStudy (anonymous):

IS IT B?

jimthompson5910 (jim_thompson5910):

the given slope is ______

OpenStudy (anonymous):

-2 AND 5

jimthompson5910 (jim_thompson5910):

that's the given point

OpenStudy (anonymous):

SO ITS A THEN

jimthompson5910 (jim_thompson5910):

the x and y coordinates more specifically

jimthompson5910 (jim_thompson5910):

how did you get A

OpenStudy (anonymous):

because therews a +5 and a -2

jimthompson5910 (jim_thompson5910):

keep in mind that you have x-x1 x1 is -2 so... x-x1 = x-(-2) = x+2

jimthompson5910 (jim_thompson5910):

so it's like the signs flip

OpenStudy (anonymous):

so i change it?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

ok i trhink its B

jimthompson5910 (jim_thompson5910):

y - y1 = m(x - x1) y - y1 = -1(x - x1) y - 5 = -1(x - (-2)) y - 5 = -1(x + 2)

jimthompson5910 (jim_thompson5910):

you somehow flipped x and y

jimthompson5910 (jim_thompson5910):

in (-2, 5) x comes first, so x = -2, y = 5

OpenStudy (anonymous):

:)

jimthompson5910 (jim_thompson5910):

think of it alphabetically